A house uses a heat pump to maintain the indoor temperature at 22 C on a cold winter day when the outside temperature is -10 C. On this day, the house receives heat from the heat pump at a rate of 24 kW. Is it possible for this heat pump to operate at 2.5 kW? Justify your answer.
Answer #1


SOL TH=22°C 27273+22=295 K Tc=2-10° c = 263K dier24 kW Win=2.5 kW 7)C:0:P of carnot (reversible) Heat Pump: (MB) Win Tie - 2